16.) Use Cramers Rule to solve for y: 2x - 3y + 4z = 1 + 6z=0 3x - 2y = 5 .
Answer #1

22-34+42=1 (toyt +04+ 62=0 yo kindly Thumbs up! If it really helped You 3x-2y+02=5.. 2y+02 2 -3 0 4 6. is coefficient matrix